# Antenna Coding Empowered by Pixel Antennas

### v1

@@ III-C1 Partition Optimization
$\mathcal{H}_{m}^{\left(i\right)}$ =$=\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{m}^{\left(i-1\right)}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}\right.$
$\left.\qquad\hfill\geq\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{m^{\prime}}^{\left(i-1\right)}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall l,\forall m^{\prime}\neq m\right\},\forall m.$ (22)
where $\mathbf{b}_{m}^{\left(i-1\right)}$ is the antenna coder optimized at iteration $i-1$.
@@ III-C2 Antenna Coder Optimization
$\mathbf{b}_{m}^{\left(i\right)}$ $=\underset{\left[\mathbf{b}_{m}\right]_{q}\in\left\{0,1\right\},\forall m,q}{\textrm{argmax}}\sum_{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\in\mathcal{H}_{m}^{\left(i\right)}}\left|\mathbf{e}_{\textrm{R}}^{H}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall m,$ (23)
@@ III-D Performance Analysis
$\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{oc}}\mathbf{i}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{U}\mathbf{S}\mathbf{V}^{H}\mathbf{i}\left(\mathbf{b}_{\textrm{R}}\right).$ (25)
@@
$h\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{i}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{V}\mathbf{S}^{H}\mathbf{U}^{H}\mathbf{H}_{\textrm{v}}\mathbf{e}_{\textrm{T}}=\mathbf{w}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\widetilde{\mathbf{h}},$ (26)
@@
$\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{S}\mathbf{V}^{H}\mathbf{i}\left(\mathbf{b}_{\textrm{R}}\right),$ (27)
@@
$\widetilde{\mathbf{h}}=\mathbf{U}^{H}\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}},$ (28)
where $[\widetilde{\mathbf{h}}]_{i}$ $\forall i$ are i.i.d. random variables following the complex Gaussian distribution $\mathcal{CN}\left(0,1\right)$ because of the $R$ orthogonal radiation patterns collected in $\mathbf{U}$.
@@ IV-A Beamspace Channel Model
$\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{R}}^{H}\left(\mathbf{B}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{E}_{\textrm{T}}\left(\mathbf{B}_{\textrm{T}}\right)$ (32)
@@ IV-C1 Antenna Coding Design with Uniform Power Allocation
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}$ $\log_{2}\left|\mathbf{I}+\frac{P}{\sigma^{2}N_{\textrm{T}}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (39)
@@ IV-C2 Joint Antenna Coding and Waterfilling Design
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}$ $\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right)$ (45)
@@ IV-D Antenna Coding Design with Codebook
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}$ $\log_{2}\left|\mathbf{I}+\frac{P}{\sigma^{2}N_{\textrm{T}}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (48)
@@
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}$ $\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right)$ (51)
@@ V-A Simulation Setup and Pixel Antenna Design
We consider two examples of pixel antennas operating at 2.4 GHz where the wavelength is $\lambda=125$ mm. The two pixel antennas are designed based on discretizing the radiating surface of conventional microstrip patch antennas [^3] Herein we consider the pixelated microstrip patch antenna for illustration. However, it should be noted that pixel antenna is general and can be designed based on discretizing the continuous radiating surface of different types of antennas such as monopole antenna, not restricted to microstrip patch antenna. into a grid of pixels, but having different physical apertures, $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$. The geometry and dimension for the two pixel antennas are illustrated in Fig. 6. For both pixel antennas, there are in total 40 ports, including 1 antenna port and $Q=39$ pixel ports, and the pixel size is $5\textrm{mm}\times 5\textrm{mm}$ and $11\textrm{mm}\times 11\textrm{mm}$, respectively. The pixels and ground plane are printed on a 1.524-mm thick Rogers 4003C substrate with a permittivity of 3.55 and a loss tangent of 0.0027. We use CST studio suite, a full-wave EM solver, to simulate the $(Q+1)$-port circuit network of pixel antennas to achieve the impedance matrix $\mathbf{Z}\in\mathbb{C}^{(Q+1)\times(Q+1)}$ as well as the open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}\in\mathbb{C}^{2K\times(Q+1)}$.
@@ V-B SISO Pixel Antenna System Performance
We first evaluate the performance of SISO pixel antenna system. Utilizing Monte Carlo method, we generate multiple realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ and for each realization, we use the SEBO algorithm [37], with block size of $J=10$, to solve the problem (12)-(13) so as to maximize the channel gain $\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}$ with perfect CSI. In addition, we also maximize the channel gain with the codebook design for antenna coding introduced in Section III-C. The average channel gain of the SISO pixel antenna system optimized with perfect CSI and with the codebook is shown in Fig. 8. The two pixel antennas with different physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are considered and the conventional SISO system is also compared as the benchmark. From Fig. 8, we can make three observations.
@@
Thirdly, for the SISO pixel antenna system, increasing the physical aperture is beneficial to enhance channel gain. For example, when optimized with perfect CSI, the channel gain with physical aperture $0.5\lambda\times 0.5\lambda$ is 35% higher than that with $0.25\lambda\times 0.25\lambda$. This is because pixel antennas with larger physical aperture can provide more orthogonal radiation patterns, i.e. larger EADoF. Specifically, the EADoF for the pixel antenna with physical aperture $0.25\lambda\times 0.25\lambda$ is 5 while the EADoF with $0.5\lambda\times 0.5\lambda$ is 7. Therefore, pixel antennas with larger physical aperture achieve a higher upper bound for the average channel gain (31), resulting in a higher average channel gain. Moreover, it is worth noting that the upper bound of the average channel gain of SISO pixel antenna system (31) is generally hard to achieve. For example, for the physical aperture $0.25\lambda\times 0.25\lambda$, the average channel gain optimized with perfect CSI is 4 while the upper bound is 5. This is because it is not guaranteed to find the antenna coder $\mathbf{b}_{\textrm{R}}$ such that the pixel antenna current $\mathbf{i}\left(\mathbf{b}_{\textrm{R}}\right)$ and the $\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)$ given by (27) satisfy the condition of maximum ratio combiner (30).
@@ V-C MIMO Pixel Antenna System Performance
We next evaluate the performance of MIMO pixel antenna systems. Similarly, we utilize the Monte Carlo method to generate multiple realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$. For each realization, we optimize the antenna coding design with uniform power allocation by solving the problem (39)-(41) and also jointly optimize the antenna coding and waterfilling design by solving the problem (45)-(47), for maximizing the channel capacity of MIMO pixel antenna system with perfect CSI. The SEBO algorithm [37], with block size $J=10$, is utilized to solve the two problems. The channel capacity of the MIMO pixel antenna system with uniform power allocation (UP) and waterfilling (WF) optimized with perfect CSI is shown in Fig. 9. The channel capacity is averaged over multiple channel realizations. The two pixel antennas with different physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are considered and the conventional MIMO system with UP and WF are compared as the benchmarks. Different numbers of transmit and receive antennas $N_{\textrm{T}}$ and $N_{\textrm{R}}$ are also considered for the MIMO pixel antenna system and conventional MIMO system. From Fig. 9, we can make four observations.

### v2

@@ III-C1 Partition Optimization
$\mathcal{H}_{m}^{\left(i\right)}$ $=\left\{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mid\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}^{\left(i-1\right)}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2}\right$.
$\left.\qquad\hfill\geq\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m^{\prime}}^{\left(i-1\right)}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall l,\forall m^{\prime}\neq m\right\},\forall m.$ (35)
where $\mathbf{b}_{m}^{\left(i-1\right)}$ is the antenna coder optimized at iteration i1i-1.
@@ III-C2 Antenna Coder Optimization
$\mathbf{b}_{m}^{\left(i\right)}$ $=\underset{\left[\mathbf{b}_{m}\right]_{q}\in\left\{0,1\right\},\forall m,q}{\textrm{argmax}}\sum_{\mathbf{H}_{\textrm{V}}^{\left[l\right]}\in\mathcal{H}_{m}^{\left(i\right)}}\left|\mathbf{e}_{\textrm{R}}^{T}\left(\mathbf{b}_{m}\right)\mathbf{H}_{\textrm{V}}^{\left[l\right]}\mathbf{e}_{\textrm{T}}\right|^{2},\forall m,$ (36)
@@ III-D Performance Analysis
$\mathbf{e}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{oc}}\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{U}\mathbf{S}\mathbf{V}^{H}\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right).$ (38)
@@
$h\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{i}_{\textrm{R}}^{T}\left(\mathbf{b}_{\textrm{R}}\right)\mathbf{V}^{*}\mathbf{S}\mathbf{U}^{T}\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}}=\mathbf{w}^{H}\left(\mathbf{b}_{\textrm{R}}\right)\widetilde{\mathbf{h}}$, (39)
@@
$\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)=\mathbf{S}\mathbf{V}^{T}\mathbf{i}_{\textrm{R}}^{*}\left(\mathbf{b}_{\textrm{R}}\right),$ (40)
@@
$\widetilde{\mathbf{h}}=\mathbf{U}^{T}\mathbf{H}_{\textrm{V}}\mathbf{e}_{\textrm{T}},$ (41)
where $\widetilde{\mathbf{h}}]_{i}$ $\forall i$ are i.i.d. random variables following the complex Gaussian distribution $\mathcal{CN}\left(0,1\right)$ because of the $R$ orthogonal radiation patterns collected in $\mathbf{U}$.
@@ IV-A Beamspace Channel Model
$\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)=\mathbf{E}_{\textrm{R}}^{T}\left(\mathbf{B}_{\textrm{R}}\right)\mathbf{H}_{\textrm{V}}\mathbf{E}_{\textrm{T}}\left(\mathbf{B}_{\textrm{T}}\right)$ (45)
@@ IV-C1 Antenna Coding Design with Uniform Power Allocation
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\log_{2}\left|\mathbf{I}+\frac{P}{\sigma^{2}N_{\textrm{T}}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (52)
@@ IV-C2 Joint Antenna Coding and Waterfilling Design
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right)$ (58)
@@ IV-D Antenna Coding Design with Codebook
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\log_{2}\left|\mathbf{I}+\frac{P}{\sigma^{2}N_{\textrm{T}}}\mathbf{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\mathbf{H}^{H}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\right|$ (61)
@@
$\underset{\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}}{\text{max}}\$$\sum_{i=1}^{N_{\textrm{min}}}\log_{2}\left(1+\frac{P_{i}^{\star}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)\lambda_{i}\left(\mathbf{B}_{\textrm{T}},\mathbf{B}_{\textrm{R}}\right)}{\sigma^{2}}\right)$ (64)
@@ V-A Simulation Setup and Pixel Antenna Design
We consider two examples of pixel antennas operating at 2.4 GHz where the wavelength is $\lambda=125$ mm. The two pixel antennas are designed based on discretizing the radiating patch surface of conventional microstrip patch antennas [^5] Herein we consider the pixelated microstrip patch antenna for illustration. However, it should be noted that pixel antenna is general and can be designed based on discretizing the continuous radiating surface of different types of antennas such as monopole antenna, not restricted to microstrip patch antenna. into a grid of pixels, but having different physical apertures, $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$. The geometry and dimension for the two pixel antennas are illustrated in Fig. 6 and summarized in Table I. For both pixel antennas, there are in total 40 ports, including 1 antenna port and $Q=39$ pixel ports, and $\times5$ pixels with pixel size being $5\textrm{mm}\times 5\textrm{mm}$ and $11\textrm{mm}\times 11\textrm{mm}$, respectively. The pixelized patch and ground plane are printed on the Rogers 4003C substrate, where the thickness is 1.524 mm, the permittivity is 3.55, and the loss tangent is 0.0027. We use CST studio suite, a full-wave EM solver, to simulate the $(Q+1)$-port circuit network of pixel antennas to achieve the impedance matrix $\mathbf{Z}\in\mathbb{C}^{(Q+1)\times(Q+1)}$ as well as the open-circuit radiation pattern matrix $\mathbf{E}_{\textrm{oc}}\in\mathbb{C}^{2K\times(Q+1)}$.
@@
TABLE I: Geometry and Dimensions for Pixel Antennas
Physical aperture $0.25\lambda\times 0.25\lambda$ $0.5\lambda\times 0.5\lambda$
Size of ground $\times31.25mm$ $\times62.5mm$
Size of patch $\times29mm$ $\times59mm$
Number of pixels $\times5$ $\times5$
Size of pixel $\times5mm$ $\times11mm$
Gap between pixels 1mm 1mm
Patch and ground spacing 12.5mm 12.5mm
@@ V-B SISO Pixel Antenna System Performance
We first evaluate the performance of SISO pixel antenna system. Utilizing Monte Carlo method, we generate multiple realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$ and for each realization, we use the SEBO algorithm [44], with block size of $J=10$ (which achieves a good optimization performance with an acceptable complexity), to solve the problem (25)-(26) so as to maximize the channel gain $\left|h\left(\mathbf{b}_{\textrm{R}}\right)\right|^{2}$ with perfect CSI. In addition, we also maximize the channel gain with the codebook design for antenna coding introduced in Section III-C. The number of virtual channel realizations in the training set is $L=15000$. The antenna coders in the codebook design with size [^6] Due to limited space, it would be messy to show the antenna coders in the codebook with size larger than 8, so we omit them. of 2, 4, and 8 are listed in Table II. The number of realizations in Monte Carlo method is 100, which is enough to obtain an accurate average performance. The average channel gain of the SISO pixel antenna system optimized with perfect CSI and with the codebook is shown in Fig. 8. The two pixel antennas with different physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are considered and the conventional SISO system is also compared as the benchmark. From Fig. 8, we can make three observations.
TABLE II: Antenna Coders in the Codebook with Size of 2, 4, 8
Physical aperture $0.25\lambda\times 0.25\lambda$ Physical aperture $0.5\lambda\times 0.5\lambda$
Codebook 000110100000011101101011000001101001101 010000110010110101101011001010011001000
size = 2 101010111110110111010001001111101010101 100010101010010110110011100110000001101
101001011100010111110010000110110001000 001010110001100101101100100100000101010
Codebook 110111011110010100110110111000101010110 111000011111110101000001000110100101111
size = 4 110100000010001000110111110000011011110 100101111010010100101000011010001101110
010110001001000101111011010111111101110 100101000010001001000010000101111100101
000000111110010010110111101001011101101 111111010000001000101111110100110110000
000111000011011101111000110111010001011 000001101000100110101010101100001000011
010011110101110101100101111011100100001 111111000111000100011000100000110001011
Codebook 001011100000010001110110100110110101100 010000110101010111010100100001001100000
size = 8 010111010110011100011011111011111101010 100100100010000000010110000101111000011
000100101101010000110110100110000101001 010111110100011100001111100100010100111
001010011111100011010111101101100000000 000101001111111101010101110100111100011
100001010010000011111000110010010111110 110000011011011000100001010010100001001
@@
Thirdly, for the SISO pixel antenna system, increasing the physical aperture is beneficial to enhance channel gain. For example, when optimized with perfect CSI, the channel gain with physical aperture $0.5\lambda\times 0.5\lambda$ is 35% higher than that with $0.25\lambda\times 0.25\lambda$. This is because pixel antennas with larger physical aperture can provide more orthogonal radiation patterns, i.e. larger EADoF. Specifically, the EADoF for the pixel antenna with physical aperture $0.25\lambda\times 0.25\lambda$ is 5 while the EADoF with $0.5\lambda\times 0.5\lambda$ is 7. Therefore, pixel antennas with larger physical aperture achieve a higher upper bound for the average channel gain (44), resulting in a higher average channel gain. Moreover, it is worth noting that the upper bound of the average channel gain of SISO pixel antenna system (44) is generally hard to achieve. For example, for the physical aperture $0.25\lambda\times 0.25\lambda$, the average channel gain optimized with perfect CSI is 4 while the upper bound is 5. This is because it is not guaranteed to find the antenna coder $\mathbf{b}_{\textrm{R}}$ such that the pixel antenna current $\mathbf{i}_{\textrm{R}}\left(\mathbf{b}_{\textrm{R}}\right)$ and the $\mathbf{w}\left(\mathbf{b}_{\textrm{R}}\right)$ given by (40) satisfy the condition of maximum ratio combiner (43).
@@ V-C MIMO Pixel Antenna System Performance
We next evaluate the performance of MIMO pixel antenna systems. Similarly, we utilize the Monte Carlo method to generate multiple realizations for the virtual channel matrix $\mathbf{H}_{\textrm{V}}$. For each realization, we optimize the antenna coding design with uniform power allocation by solving the problem (52)-(54) and also jointly optimize the antenna coding and waterfilling design by solving the problem (58)-(60), for maximizing the channel capacity of MIMO pixel antenna system with perfect CSI. The SEBO algorithm [44], with block size $J=10$, is utilized to solve the two problems. The channel capacity of the MIMO pixel antenna system with uniform power allocation (UP) and waterfilling (WF) optimized with perfect CSI is shown in Fig. 9. The channel capacity is averaged over multiple channel realizations. The two pixel antennas with different physical apertures $0.25\lambda\times 0.25\lambda$ and $0.5\lambda\times 0.5\lambda$ are considered and the conventional MIMO system with UP and WF are compared. Different numbers of transmit and receive antennas $N_{\textrm{T}}$ and $N_{\textrm{R}}$ are also considered for the MIMO pixel antenna system and conventional MIMO system. From Fig. 9, we can make four observations.